### v2

# Note on Klein-Nishina effect in strong-field QED: the case of nonlinear Compton scattering

Abstract
Suitably normalized differential probabilities of one-photon emission in external electromagnetic fields are compared to quantify the transit of nonlinear Compton scattering to linear Compton scattering, described by the Klein-Nishina formula, and to constant crossed field treatment. The known Klein-Nishina suppression at large energies is further enforced by increasing field intensity. In view of the Ritus-Narozhny conjecture, we demonstrate that different paths in the field intensity vs. energy plane towards large values of the quantum non-linearity parameter $\chi$ facilitate significantly different asymptotic dependencies, both in the Klein-Nishina regime and the constant crossed field regime and in between.
strong-field QED: nonlinear Compton scattering, linear Compton scattering, Klein-Nishina effect, constant crossed field, Ritus-Narozhny conjecture
I Introduction
The Klein-Nishina effect means a dropping cross section of the linear Compton scattering Klein:1929lcc from Thomson’s value at small energies towards high energies. In astrophysics, this dropping reduces the energy loss of ultra-relativistic electrons and positrons Fang:2020dmi ; Schlickeiser:2009qq and the inverse Compton scattering rates with ultra-high energy photons, most notable for energies in the order of PeV LHAASO:2023gne ; LHAASO:2021gok . Additionally, in certain sites, ultra-intense electromagnetic background fields are present and enable or influence elementary charged-particle–photon interactions. Pulsars in general and magnetars in particular represent such sites. In earth-bound laboratories similarly extreme conditions can be achieved, hence linking to or exploring analog astrophysics scenarios. Beam-beam interactions in accelerators Yakimenko:2018kih or beam-laser interactions Abramowicz:2021zja ; Salgado:2021fgt ; Storey:2023wlu or high-intensity laser-electron (plasma Brodin:2022dkd ; Seipt:2023bcw ; Zhang:2020lxl ) experiments Poder:2017dpw ; Cole:2017zca or laser-laser interactions Blinne:2018nbd ; Gies:2017ezf ; Gies:2017ygp are to be mentioned in that context.
A special branch of phenomenology is tied to strong-field Quantum Electro Dynamics (sfQED) (see Fedotov:2022ely for a recent survey with an extensive citation list and HernandezAcosta:2023msl for a concise formulation in momentum space) aiming at describing the fundamental electromagnetic interactions in or at a background field. The latter one could be of electromagnetic (or electro-weak) nature or being gravity or something - hypothetically - beyond the Standard Model of Particle Physics. We confine ourselves to electromagnetic backgrounds, i.e. sfQED.111 Many entries on the topic are cited in Fedotov:2022ely . Further surveys are provided by Gonoskov:2021hwf ; Titov:2015pre ; DiPiazza:2011tq ; Kaminski:2009wwd .
A straightforward question is then whether a noticeable background field has a significant impact on elementary processes. We narrow here the problem to the impact of an electromagnetic field on the photon emission off electrons and consider specifically Compton scattering.222 A few entries for Compton as one-photon emission are Ritus85 ; Lotstedt:2009zz ; Mackenroth:2010jr ; Hartin:2011vr ; Heinzl:2009nd ; Seipt:2010ya ; Seipt:2011zz ; Seipt:2011dx ; Seipt:2013taa ; Seipt:2016rtk ; Titov:2015tdz ; Dinu:2018efz ; Ilderton:2018nws ; Ilderton:2020dhs ; Nielsen:2021nuo ; Gelfer:2022nqy ; Kampfer:2020cbx ; Acosta:2021iyu ; Titov:2023yev ; Angioi:2016vir ; Blackburn:2018sfn ; for two-photon Compton (two-vertex diagram) see, e.g., Herrmann:1971jc ; Guccione-Gush:1975juz ; Akhiezer:1985 ; Ehlotzky:1989 ; Mackenroth:2012rb ; Seipt:2012tn ; Dinu:2018efz ; Seipt:2013hda , and higher orders in Lotstedt:2012zz ; Lotstedt:2013uya ; Dinu:2019pau .
One aspect concerns the Klein-Nishina effect in a strong background field. Another issue is related to the Ritus-Narozhny conjecture, arguing that the account of the background field within the Furry picture is restricted to not too intense fields Ritus:1972ky ; Fedotov:2016afw ; Heinzl:2021mji ; Mironov:2020gbi . In doing so we resort to fairly known formalism as a vehicle to elucidate, numerically, links between established descriptions of linear Compton scattering, nonlinear Compton scattering, and photon emission in a constant crossed field. This subject is also of interest for approximations implemented in simulation codes - both for astrophysics applications and laser physics in particular (cf. DiPiazza:2018bfu ; DiPiazza:2017raw ; Gelfer:2022nqy ; Nielsen:2021nuo ; Blackburn:2021cuq ; Dinu:2019pau ; Ilderton:2018nws ; Heinzl:2020ynb ).
Our note is organized as follows. In section II, we recall formulas describing the normalized probabilities of linear and nonlinear Compton scattering as well as photon emission in a constant crossed field. Subsequently, the dependence on a Lorentz invariant quantity characterizing the emitted photon (Section III), the energy dependence (Section IV) and external field intensity dependence (Section V) are analyzed. Section VI demonstrates how various paths approach a large value of the so-called quantum non-linearity parameter $\ chi$ leaves a strong impact on the asymptotic. After a brief discussion in Section VII, we conclude in Section VIII. A few relevant numbers are listed in Appendix A, and Appendix B recaps relations of plane wave and constant crossed field quantities. Appendix C presents master curves of the harmonics structures. Appendix D relates the invariant differential cross section and the photon-energy and photon-angular differential cross sections.
II Exclusive one-photon emission: recollection of linear Compton vs. nonlinear Compton vs. constant crossed field
The Klein-Nishina (KN) effect means a dropping cross section of the linear Compton scattering Klein:1929lcc from Thomson’s value $\sigma^{\mathrm{Thomson}}=\frac{8\pi}{3}r_{e}^{2}$ at small energies, $p\cdot k/m^{2}=(s-m^{2})/m^{2}\to 0$, towards high energies:333 We employ units with $\hbar=c=1$ and note the electron radius of $r_{e}=\alpha/m$ with Sommerfeld’s fine-structure constant $\alpha\approx 1/137$ (ignoring for the sake of clarity the running coupling) and electron mass $m$; the electron’s and photon’s $in$ four-momenta are $p$ and $k$, respectively (for $out$-states, the four-momenta receive primes). The Mandelstam variable $s=(p+k)^{2}$ refers to the notion of “energy”.
$\sigma^{\mathrm{KN}}$ $=\frac{2\pi r_{e}^{2}}{u_{KN}}\left[\left(1-\frac{4}{u_{\mathrm{KN}}}-\frac{8}{u_{\mathrm{KN}}^{2}}\right)\log(1+u_{\mathrm{KN}})+\frac{1}{2}+\frac{8}{u_{\mathrm{KN}}}-\frac{1}{2(1+u_{\mathrm{KN}})^{2}}\right]$ (1)
=$=\left\{\begin{array}[]{ll}\frac{8\pi r_{e}^{2}}{3}(1-u_{\mathrm{KN}})&\mathrm{for}\,u_{\mathrm{KN}}\ll 1\\ \pi r_{e}^{2}\frac{1}{u_{\mathrm{KN}}}\left(2\log u_{\mathrm{KN}}+1\right)&\mathrm{for}\,u_{\mathrm{KN}}\gg 1\end{array}\right.,$ (4)
where $u_{\mathrm{KN}}:=2p\cdot k/m^{2}$, see empty squares in Fig. 1-left. At the origin of such behavior is the differential cross section LLIV
$\frac{d\sigma^{\mathrm{KN}}}{du}$ $=\frac{2\pi\alpha^{2}}{p\cdot k}\frac{1}{(1+u)^{2}}\left(1+\underline{\frac{u^{2}}{2(1+u)}}-2\frac{u}{u_{KN}}\left[1-\frac{u}{u_{\mathrm{KN}}}\right]\right)$ (5)
$=\frac{2\pi\alpha^{2}}{p\cdot k}\times\left\{\begin{array}[]{ll}1\,-2u(1+u_{\mathrm{KN}}^{-1})+u^{2}\left(\frac{7}{2}\frac{4}{u_{\mathrm{KN}}}+\frac{2}{u_{\mathrm{KN}}^{2}}\right)+\mathcal{O}(u^{3})&\mathrm{for}\,u\to 0\\ \frac{2+2u_{\mathrm{KN}}+u_{\mathrm{KN}}^{2}}{2(1+u_{\mathrm{KN}})^{3}}\stackrel{{\scriptstyle u_{\mathrm{KN}}\gg 1}}{{\longrightarrow}}\frac{1}{4}\frac{m^{2}}{p\cdot k}&\mathrm{for}\,u=u_{\mathrm{KN}}\\ \end{array}\right.,$ (8)
where $u\in[0,u_{KN}]$ denotes the Ritus invariant $u=\frac{k\cdot k^{\prime}}{k\cdot p^{\prime}}$ Ritus85 which is related, via energy-momentum balance $p+k=p^{\prime}+k^{\prime}$, to the lightfront momentum fraction of the $out$-photon $\mathfrak{s}=\frac{k\cdot k^{\prime}}{k\cdot p}$ via $\mathfrak{s}=\frac{u}{1+u}$. The underlined term in Eq. (5) gives rise to the logarithmic high-energy behavior $\sigma^{\mathrm{KN}}\propto(p\cdot k/m^{2})^{-1}\log(p\cdot k/m^{2})$. With increasing energy, the differential cross section drops at small values of $u$ as $\propto(p\cdot k/m^{2})^{-1}$ which turns, for $u>1$, into $\propto(p\cdot k/m^{2})^{-2}$, see Fig. 1-right. Quite qualitatively, one can argue that, with increasing energy, more and more exclusive and inclusive channels open with the consequence that one particular exclusive channel receives less probability of keeping the unitarity of the S matrix. In the case of linear Compton scattering, often the electron recoils are said to reduce the $out$ photon energy.
[…]
The above (differential) cross section is for the leading-order tree-level term of exclusive $2\to 2$ scattering $e^{-}+\gamma\to{e^{-}}^{\prime}+\gamma^{\prime}$ within perturbative QED, organized in powers of $\alpha$ (for higher orders, cf. Lee:2021iid ). It is known LLIV ; Ritus85 to emerge from nonlinear Compton scattering, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, as leading-order weak-field term, where $e_{A}^{-}$ is for the electron wave function in an external, classical background field $A$ which may refer to a monochromatic laser field in a plane-wave model as the so-called infinite plane-wave approximation (IPA). Having this relation in mind, one can define a suitably normalized probability (henceforth dubbed ”cross section”) which makes this relation explicit:444 See Bragin:2020akq for the definition of “local cross sections” for processes in an external field. Instead of using wave packages, we employ a circularly polarized, monochromatic plane-wave field as a proxy of a mono-energetic photon beam under special circumstances. The intensity parameter is $a_{0}$.
$\frac{d\sigma^{\mathrm{IPA}}}{du}$ $=\frac{2\pi\alpha^{2}}{p\cdot k}\frac{1}{(1+u)^{2}}\sum_{n=1}^{N\to\infty}\Theta(u_{n}-u)\mathcal{F}_{n}(z_{n}),$ (9)
$\mathcal{F}_{n}=-\frac{2}{a_{0}^{2}}J_{n}^{2}(z_{n})+A\left[J_{n+1}^{2}(z_{n})+J_{n-1}^{2}(z_{n})-2J_{n}^{2}(z_{n})\right]$ (10)
$z_{n}=\frac{2na_{0}}{\sqrt{1+a_{0}^{2}}}\sqrt{\frac{u}{u_{n}}\left(1-\frac{u}{u_{n}}\right)}$ (11)
$u_{n}=\frac{2n}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}\quad\to n\geq n_{min}\equiv\frac{1}{2}u(1+a_{0}^{2})\frac{m^{2}}{p\cdot k}$ (12)
$A=1+\frac{u^{2}}{2(1+u)},$ (13)
see Acosta:2021iyu ; Kampfer:2020cbx and Ritus85 for the basics. The classical field strength parameter characterizes the background field $a_{0}$, and the meaning of the variable $u$ is as above. The differential cross section depends separately on two variables: the energy variable $p\cdot k/m^{2}$ and intensity parameter $a_{0}$, which can be combined with the so-called “quantum non-linearity” parameter $\chi=a_{0}\,p\cdot k/m^{2}$ and a second one. The total cross section555 The subsequent numerical evaluations uncover a large range of numerical values of the involved parameters. Therefore, the summation of a large number of harmonics, $N=\mathcal{O}(10^{9})$ in Eq. (9), is often required. (Based on properties of the Bessel functions of large arguments and large indices, harmonics up $N=\mathcal{O}(a_{0}^{3})$ should be accounted for.) In accomplishing that goal, we employ the series expansion method referred to as Richardson extrapolation Sidi:2003bnm , which drastically improves the rate of conversions. $\sigma^{\mathrm{IPA}}$ is compared with $\sigma^{\mathrm{KN}}$ in Fig. 1-left. In addition to the KN effect, the external field causes a further suppression, increasing with $a_{0}$. In line with the above-quoted argument, one can expect that, according to unitarity. In other words, the exclusive decay of an electron state, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, quantified by the total cross section, becomes more and more reduced with increasing environmental field $A$. The degree of suppression depends on the energy, at a given field strength $A$. This is our executive summary, which is put subsequently into various perspectives in a differential manner. As far as we know, this knowledge appears to be relatively uncommon.
Sometimes, it is claimed in the literature that large values of $\chi$ would refer to high intensities, and the differential probability behind Eq. (9) turns in the expression for one-photon emission in a constant crossed field (CCF). Transferred to “cross section”, linked in an ad hoc manner to Eq. (9), with some seemingly abuse of relations of $\chi$ and $a_{0}$ and energy (see however King:2014wfa for a rationale and discussion in Appendix B), one has  [^6] The locally constant field approximation (LCFA) uses the rate $\frac{dR_{\mathrm{LCFA}}}{d\mathfrak{s}}=\frac{d\sigma^{\mathrm{CCF}}_{a_{0}\to a_{0}g(\phi)}}{d\mathfrak{s}}\frac{a_{0}^{2}g(\phi)^{2}}{4\pi\alpha m^{2}}$, related to CCF Eqs. (14 - 16), where $\mathfrak{s}=u/(1+u)$ is the above mentioned lightfront momentum transfer fraction, see Fedotov:2022ely for details. The quantity $g(\phi)$ is the pulse envelope describing the pulse evolution as a function of the invariant phase $\phi$. Analogously, the locally monochromatic approximation (LMA) is related to IPA via $\frac{dR_{\mathrm{LMA}}}{d\mathfrak{s}}=\frac{d\sigma^{\mathrm{IPA}}_{a_{0}\to a_{0}g(\phi)}}{d\mathfrak{s}}\frac{a_{0}^{2}g(\phi)^{2}}{4\pi\alpha m^{2}}$, where in IPA Eqs. (9 - 13) the replacement $a_{0}\to a_{0}g(\phi)$ also applies and $u=s/(1-s)$ is to be used. (Mimicking pulses by $a_{0}\to a_{0}g(\phi)$ has been traded, according to Heinzl:2020ynb for the first time, in Titov:2019kdk in an ad hoc manner.) Both LCFA and LMA are often used approximation schemes, implemented in many simulation codes for laser-particle/plasma interactions. LMA is shown in Heinzl:2020ynb to approach LCFA in the strong-field limit. The rates $R_{\mathrm{LCFA}}$ and $R_{\mathrm{LMA}}$, defined by Eqs. (84) and (98) in Fedotov:2022ely , require the $\mathfrak{s}$ and $\phi$ integrations, see also Eqs. (1, 2) in Ilderton:2018nws . For more details on LCFA and LMA and improvements, see DiPiazza:2018bfu ; DiPiazza:2017raw ; Ilderton:2018nws .
$\frac{d\sigma^{\mathrm{CCF}}}{du}$ $=-\frac{4\pi\alpha^{2}}{m^{2}a_{0}}\frac{1}{\chi}\frac{1}{(1+u)^{2}}\mathcal{F}_{\mathrm{CCF}}(\hat{z}),$ (14)
[…]
$\hat{z}=\left(\frac{u}{\chi}\right)^{2/3},\quad\chi=a_{0}\frac{p\cdot k}{m^{2}}\mathrm{~{}is~{}independent~{}of~{}}\omega,$ (16)
which in turn obeys the asymptotic Ritus85
$\frac{d\sigma_{\mathrm{large}-u}^{\mathrm{CCF}}}{du}=\frac{2\sqrt{\pi}\alpha^{2}}{m^{2}a_{0}}\chi^{-1/2}u^{-3/2}\exp\left\{-\frac{2u}{3\chi}\right\},\quad\mathrm{for~{}}u\gg 1,\chi.$ (17)
[…]
Given that specific framework, we quantify the relations of the above-quoted cross sections on a differential level and point out the detailed conditions for $d\sigma^{\mathrm{IPA}}/du\to d\sigma^{\mathrm{CCF}}/du$ and $d\sigma^{\mathrm{IPA}}/du\to d\sigma^{\mathrm{KN}}/du$. We also show that approaching large values of $\chi$ along different paths in the $a_{0}$-$p\cdot k/m^{2}$ plane facilitates different asymptotic.
III The $\mathbf{u}$ dependence
Figure 2 exhibits the differential cross section $d\sigma^{\mathrm{IPA}}/du$ as a function of $u$ for various values of $a_{0}=10^{0}\cdots 10^{3}$. Squares display the KN cross section, while the CCF cross section is marked by dots. The left panel is for $p\cdot k/m^{2}=10^{-1}$, while the right panel has $p\cdot k/m^{2}=10^{1}$. With increasing values of $a_{0}$ the “spectra” are shifted to the left, as also do the harmonic structures. The behavior in the IR region, i.e. at small $u$, is determined by
$\frac{d\sigma^{\mathrm{IPA}}}{du}\bigg{|}_{u\to 0}\approx\frac{2\pi\alpha^{2}}{p\cdot k}\left\{1-u\left[2+\frac{1+2a_{0}^{2}}{p\cdot k/m^{2}}\right]\right\}\approx\frac{d\sigma^{\mathrm{KN}}}{du}\bigg{|}_{\mathrm{small~{}}u}-\frac{2\pi\alpha^{2}}{p\cdot k}\,u\,\frac{2a_{0}^{2}}{p\cdot k/m^{2}},$ (18)
[…]
Figure 2: Differential cross section as a function of $u$ for $p\cdot k/m^{2}=0.1$ (top panel), 1 (left bottom panel) and 10 (right bottom panel). IPA, Eq. (9), for various values of $a_{0}=1,10^{1},10^{2},10^{3}$ and $10^{4}$ (from top to bottom), is exhibited by solid curves. The squares are for KN, Eq. (5). Circles depict CCF, Eq. (14), beyond the harmonic structures. (At smaller values of $u$, CCF fails badly by a strong overshot, even diverging for $u\to 0$.) Right/below the thin dashed curves, 90% of the total cross sections is located.
As an interim summary, we mention: $d\sigma^{\mathrm{IPA}}/du\stackrel{{\scriptstyle\mathrm{small}~{}u}}{{\xrightarrow{\hskip 25.6073pt}}}d\sigma^{\mathrm{KN}}/du$, $d\sigma^{\mathrm{IPA}}/du\stackrel{{\scriptstyle\mathrm{large}~{}u}}{{\xrightarrow{\hskip 25.6073pt}}}d\sigma^{\mathrm{CCF}}/du$ when keeping fixed $a_{0}$ and $p\cdot k/m^{2}$; the notions “small”/“large” depend separately on $a_{0}$ and $p\cdot k/m^{2}$. The harmonic structure at $u=u_{1}$ may serve as demarcation via
$\frac{d\sigma^{\mathrm{IPA}}}{du}\bigg{|}_{u=u_{1}}=\frac{2\pi\alpha^{2}}{p\cdot k}\frac{1}{(1+u_{1})^{2}}\left(1+\frac{u_{1}^{2}}{2(1+u_{1})}\right)\approx\frac{d\sigma^{\mathrm{KN}}}{du}\bigg{|}_{u=0}\times\left\{\begin{array}[]{ll}1&\mathrm{for~{}}u_{1}\ll 1,\\ \frac{1+a_{0}^{2}}{4}\frac{m^{2}}{p\cdot k}&\mathrm{for~{}}u_{1}\gg 1,\\ \end{array}\right.$ (21)
where $u_{1}\stackrel{{\scriptstyle\gg}}{{\ll}}1$ corresponds to $\frac{p\cdot k}{m^{2}}\stackrel{{\scriptstyle\ll}}{{\gg}}1+a_{0}^{2}$.
IV The energy dependence
Turning now to the energy dependence we exhibit in Fig. 3 the differential IPA cross section for $u=0.1$ (left panel), $u=1$ (middle panel) and $u=10$ (right panel). For a given value of $a_{0}$, the IPA cross section matches the CCF expression at small energies, while the high-energy tails coincide with KN. The demarcation is again provided by the first harmonic structures depending on $a_{0}$ and $p\cdot k/m^{2}$ at a given value of $u$ via $u=u_{1}=\frac{2}{1+a_{0}^{2}}p\cdot k/m^{2}$. As mentioned above, besides the total cross section, the differential KN cross section also drops with increasing energy $p\cdot k/m^{2}$. On top of that dropping is the suppression of the differential IPA and CCF cross sections at small energies.
Figure 3: Differential IPA cross section (solid curves) for $u=0.1$ (top panel), 1 (left bottom panel) and 10 (right bottom panel) as a function of energy for various values of $a_{0}=1,10^{1},10^{2},10^{3}$ and $10^{4}$ (from top to bottom). The squares are for KN, and dots depict CCF beyond harmonic structures.
As an interim summary, we mention: $d\sigma^{\mathrm{IPA}}/du\stackrel{{\scriptstyle\mathrm{small}~{}p\cdot k/m^{2}}}{{\xrightarrow{\hskip 42.67912pt}}}d\sigma^{\mathrm{CCF}}/du$, $d\sigma^{\mathrm{IPA}}/du\stackrel{{\scriptstyle\mathrm{large}~{}p\cdot k/m^{2}}}{{\xrightarrow{\hskip 42.67912pt}}}d\sigma^{\mathrm{KN}}/du$ offer a different perspective when keeping fixed $u$ and $a_{0}$. Here, “small energies” mean $p\cdot k/m^{2}<\frac{1}{2}u(1+a_{0}^{2})$, while “large energies” mean $p\cdot k/m^{2}>\frac{1}{2}u(1+a_{0}^{2})$. Since this is the edge of the first harmonic, $p\cdot k/m^{2}$ needs to be somewhat smaller for CCF and somewhat bigger for KN.
V The $\mathbf{a_{0}}$ dependence
To identify further regions of dominating CCF behavior and discriminate against the KN regime, we display in Fig. 4 the differential cross section at $u=0.1$ (right panel), $u=1$ (middle panel) and $u=10$ (right panel) as contour plots over the $a_{0}$-$p\cdot k/m^{2}$ plane. The differential cross section clearly shows that even for such large values of $\chi$ as $10^{4}$ one encounters the KN regime at large energy $p\cdot k/m^{2}>10^{2}$. The KN regime is that where the contour lines are vertical, i.e. independent of $a_{0}$. In contrast, even for small values of $\chi$, e.g. $10^{0}$, the inclined contour lines at small energies, e.g. $p\cdot k/m^{2}=10^{-1}$, point to the CCF regime. The line $u_{1}=u$, at given $u$, is the demarcation, with a structure related to the lowest harmonics. (Appendix C focuses on the harmonic structures.)
Left to the dashed vertical line the KN regime terminates due to kinematic reasons. In the deep blue region, the exponential suppression (17) applies.
Figure 4: Contour plots of $d\sigma^{\mathrm{IPA}}/du$ in units of mb over the $a_{0}$-$p\cdot k/m^{2}$ pane. Lines of $\chi=const$ are also displayed (dotted). Top/left bottom/right bottom panels are for $u=0.1/1/10$.
Going, at $p\cdot k/m^{2}=const$, from the KN region to larger values of $a_{0}$ in the CCF region, one crosses the inclined contour lines, meaning a dropping $d\sigma^{\mathrm{IPA}}/du$. This, together with the $u$ dependence, is at the origin of the dropping total cross section with increasing $a_{0}$ exhibited in Fig. 1-left.
Coming back to Figs. 2 and 3, one infers that increasing values of $a_{0}$ suppress the differential cross section (see Fig. 2 for $p\cdot k/m^{2}=const$ and Fig. 3 for $u=const$) at medium-$u$, while enabling non-zero contributions beyond the KN limit, i.e. at $u>u_{\mathrm{KN}}$. The net effect is nevertheless the suppression of the total cross section, as exhibited in Fig. 1-left. That is, the Klein-Nishina effect is amplified by the external field. Both, differential and total cross section are modified. Reference Lotstedt:2008 reports on a different example, where an external field modifies the differential cross section, while the total cross section remains nearly unchanged. This is similar to the findings in Seipt:2013hda . The difference is in field-enabled vs. field-assisted processes.
As an interim summary let us mention that a more differential perspective, useful for defining “KN regime” and “CCF regime” at $u=const$, is Fig. 4: The KN regime is below the line $u_{1}(a_{0},p\cdot k/m^{2})=u$, i.e. $a_{0}^{(1)}=\sqrt{\frac{2}{u}\frac{p\cdot k}{m^{2}}-1}$, which becomes a near-vertical line $p\cdot k/m^{2}=u/2$ for $a_{0}<1$. The contour lines there are vertical, i.e. independent of $a_{0}$. The CCF regime, where contour lines are inclined due to the $a_{0}$ dependence, is above the curves $a_{0}^{(N)}=\sqrt{\frac{2N}{u}\frac{p\cdot k}{m^{2}}-1}$. In the boomerang-shaped corridor $a_{0}^{(1)}(p\cdot k)\cdots a_{0}^{(N)}(p\cdot k)$, the first $N$ harmonics can be recognized; their visible structures fade away for $N>6$, see also Appendix C. The CCF regime is apparently limited to $\chi>u/10$ at smaller values of $p\cdot k/m^{2}$ and larger $a_{0}$. Note the back bending of contour lines at $\chi\approx u$ which approach $\chi=u/10$ at $a_{0}>1$. The color code suggests a Y-shaped structure, with the deep blue region belonging to the exponential suppression, see also the left flanks in Fig. 3. (Since $u$ is prescribed and the available energy $p\cdot k/m^{2}$ in the entrance channel is too small, even high harmonics do not noticeably contribute to the differential cross section, so that the CCF region becomes V-shaped as the upper part of the Y.)
This supplements the summary of the previous section: When keeping fixed $u$ and $p\cdot k/m^{2}$, then $d\sigma^{\mathrm{IPA}}/du\stackrel{{\scriptstyle\mathrm{large}~{}a_{0}}}{{\xrightarrow{\hskip 28.45274pt}}}d\sigma^{\mathrm{CCF}}/du$, $d\sigma^{\mathrm{IPA}}/du\stackrel{{\scriptstyle\mathrm{small}~{}a_{0}}}{{\xrightarrow{\hskip 28.45274pt}}}d\sigma^{\mathrm{KN}}/du$.
The $a_{0}$-$p\cdot k/m^{2}$ plane can be divided in three regions, YKNexp.supp.CCF at $u=const$, all captured by $d\sigma^{\mathrm{IPA}}/du$ at once. Such a survey might help discuss applicability domains of LCFA and LMA which however require the $u$ integration to arrive at total rates, as pointed out in footnote 6. Our consideration suggest that, on a differential level, LMA turns also into LCFA at high intensities.
VI Many roads lead to large $\chi$
The $a_{0}$-$p\cdot k/m^{2}$ plane can be mapped out by many other coordinates, e.g. $\chi=a_{0}\,p\cdot k/m^{2}$ plus a second one, such as $a_{0}$ or $p\cdot k/m^{2}$ or (functions of) combinations thereof. An example of orthogonal coordinates is the pair $\chi,\eta^{2}$ with $\eta^{2}(a_{0},p\cdot k/m^{2})=(p\cdot k/m^{2})^{2}-a_{0}^{2}$ for $\eta^{2}\in[-\infty,\infty]$ and our $\chi(a_{0},p\cdot k/m^{2})=a_{0}\,p\cdot k/m^{2}$ for $\chi\in[0,\infty]$.
Trading CCF as proxy of $\frac{d\sigma^{IPA}}{du}$, the differential cross section can be fitted by $\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u=const}=\mathcal{A}(u)\,\chi^{-x}$ for $u\in[0.1,20]$ in the asymptotic region, i.e. at large $\chi$. Various trajectories of going to large $\chi$ in the $a_{0}$ vs. $p\cdot k/m^{2}$ plane are exhibited in Fig. 5-left, for definiteness starting at $a_{0}=1$ and $p\cdot k/m^{2}=1$. We find for a path along $a_{0}=const$ (black line): $x=1/3$, along $a_{0}=(p\cdot k/m^{2})^{\beta}$: $x=2/3$ for $\beta=1/2$ (red line), $x=5/6$ for $\beta=1$ (blue line), $x=1$ for $\beta=2$ (green line), and along $p\cdot k/m^{2}=const$ (magenta line): $x=4/3$. That is, $x$ depends strongly on the path of going to large $\chi$. For a path $a_{0}=\mathfrak{z}\left(\frac{p\cdot k}{m^{2}}\right)^{\beta}$, one gets the $\chi$ dependence $\frac{d\sigma^{\mathrm{CCF}}}{du}\propto\mathfrak{z}^{-1/(1+\beta)}\chi^{-1/3-\beta/(1+\beta)}$. The large-$\chi$ limits considered in Ilderton:2019kqp ; Podszus:2018hnz are contained as special cases.
Figure 5-right exhibits the same information in orthogonal coordinates $\eta^{2}>0$ vs. $\chi$. For $p\cdot k/m^{2}>a_{0}$, $\eta^{2}\approx const>0$ corresponds to going to larger $\chi$ at $p\cdot k/m^{2}=const$, and vice versa $\eta^{2}\approx const<0$ corresponds to going to larger $\chi$ at $a_{0}=const$ as long as $a_{0}>p\cdot k/m^{2}$. (See remark in figure caption on the relation of $\eta^{2}\lessgtr 0$ sectors.) As said above, many ways lead to large $\chi$ (and result in fairly different asymptotic regimes).
Figure 5: Left panel: The $a_{0}$-$p\cdot k/m^{2}$ plane is mapped out by curves $\chi=const$ (grey lines). A few paths of going to large values $\chi$ are depicted, labeled by the parameter $x$ determining the asymptotic $\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u=const}\propto\chi^{-x}$. Right panel: The same information but in orthogonal coordinates $\eta^{2}>0$ vs. $\chi$. Solid and dashed curves are for $a_{0}=const$ and $p\cdot k/m^{2}=const$, respectively. (The near-vertical sections are replaced by arrows.) Heavy dots depict the trajectories with $\beta=1/2$ and 2; the trajectory with $\beta=1$ is along the $\eta^{2}=0$ line (not exhibited). Mapping out the full $a_{0}$ vs. $p\cdot k/m^{2}$ plane displayed in the left panel requires the complement $\eta^{2}<0$ vs. $\chi$ (not shown since it arises from re-labeling $a_{0}\leftrightarrow p\cdot k/m^{2}$ for the coordinate lines $a_{0}=const$ and $p\cdot k/m^{2}=const$ as well as $1/2\leftrightarrow 2$ for the trajectories).
In a nutshell: along trajectories in the $a_{0}$-$p\cdot k/m^{2}$ plane, we find $d\sigma^{\mathrm{CCF}}\propto\chi^{-x}$ with $x$ strongly depending on the actually chosen path in traversing the lines $\chi=const$. In that region, where CCF is a good approximation of IPA, this finding applies equally well to $d\sigma^{\mathrm{IPA}}/du|_{u=const}$.
VII Discussion of dependencies and regimes
The here employed formulas (1 - 16) of one-photon emission in some external univariate null-fields display often structures such as $1+u$, $1+a_{0}^{2}$, $1-u/u_{n}$, or $1+u_{\mathrm{KN}}$ etc. suggesting various regimes depending on whether $1<u$ or $1>u$, $1<a_{0}^{2}$ or $1>a_{0}^{2}$, $u_{n}<1$ or $u_{n}>1$ etc. Our numerical study has the goal of identifying the corresponding regimes, beyond asymptotic behaviors, such as $u\to 0$, $a_{0}\to 0$ or $a_{0}\to\infty$, $p\cdot k/m^{2}\to\infty$ etc. That is, we quantify the triple $u$ $a_{0}$ $p\cdot k/m^{2}$ for moving in upper part of the triangle $d\sigma^{\mathrm{KN}}$$d\sigma^{\mathrm{CCF}}$$d\sigma^{\mathrm{IPA}}$ with baseline $d\sigma^{\mathrm{KN}}\leftrightarrow d\sigma^{\mathrm{CCF}}$ given by classical emission (e.g. described by photon emission along an electron trajectory in the external field via Lienard-Wiechert potential which ignores the recoil). A key for comparing IPA and KN and CCF is the proper definition of their differential cross sections, including a suitable definition of the intensity parameter $a_{0}$ in CCF. Roughly speaking, CCF applies at $F\frac{2}{u}\frac{1}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}<1$ and KN at $F^{-1}\frac{2}{u}\frac{1}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}>1$, where $F=1+\mathcal{O}(1)$ is a fudge factor aimed at accommodating the region of first harmonics. In other words, the three-dimensional parameter space has a layer from $u=F^{-1}\,\frac{2}{u}\frac{1}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}$ to $u=F\,\frac{2}{u}\frac{1}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}$; above (below) that layer, CCF (KN) approximates fairly well the invariant differential IPA cross section.
It comes as some surprise that $d\sigma^{\mathrm{IPA}}_{\mathrm{circ.\,pol.}}\to d\sigma^{\mathrm{CCF}}$, since circular polarization (circ. pol.) means a rotating electric field, i.e. one would expect a closer correspondence of $d\sigma^{\mathrm{IPA}}_{\mathrm{lin.\,pol.}}$ and $d\sigma^{\mathrm{CCF}}$ since CCF means fixed (space-time independent) electric ($E$) and magnetic ($B$) fields, which can be aligned with the linearly polarized (lin. pol.) wave fields. In Appendix B, this issue is considered, as also the meaning of $a_{0}$ w.r.t. CCF.
We prefer to consider the differential probability in the invariant Ritus variable $u$ instead of using the scattering angle or frequency $\omega^{\prime}$ of the emitted photon. For the mapping, $u\mapsto\omega^{\prime}$, cf. the discussion in Acosta:2021iyu and Appendix D.
When turning from the monochromatic external field to plane-wave pulses, the differential spectra become smoothed out, as quantified in Kampfer:2020cbx , i.e. the harmonic structures are less significant. In momentum balance equations, the electron IPA quasi-momenta are replaced by momenta, as occurring in CCF and KN, of course.
VIII Summary
Inspired by the ongoing discussion of the Ritus-Narozhny conjecture concerning the validity of the Furry picture for processes at and in external fields, we analyze numerical probabilities of linear and nonlinear Compton scattering. To establish a link to the Klein-Nishina cross section for linear Compton scattering, we chose suitably normalized probabilities as “cross sections”, mainly based on the long-known relation of the weak-field limit of the nonlinear Compton scattering matching the Klein-Nishina expression. Choosing furthermore the monochromatic plane wave model as a proxy of a mono-energetic photon beam, we show with established expressions [^9] Since we envisage a clear cut picture with well-defined precise basics, we resort to the simplest (quantum) formulas describing exclusive one-photon emission. Thus, such a compendium ignores much of the refinements elaborated during the last years, e.g. w.r.t. pulse shapes (implying bandwidth, ponderomotive and interference effects), laser focusing, and the many useful approximation schemes and simulation codes etc., including the efforts of benchmarks. By exhibiting a few relevant plots, we want to display systematically in an intriguing manner the interplay of Klein-Nishina (i.e. linear Compton) and constant-crossed fields to nonlinear Compton and identify - purely numerically - the overlap regions. For a comprehensive survey of modifications of the IPA model, see Fedotov:2022ely . In essence, the parameter space is noticeably enlarged requiring a multitude of specialized considerations of various distinctions. for the one-photon emission in pane waves and constant crossed field how and where the so-defined cross sections coincide. We do our analysis differentially with a Lorentz invariant quantity - the Ritus variable $u$ instead of frequency or/and scattering angle of the produced photon emitted by an electron moving in the external field or scattering at a photon beam. We identify those regions in the three-dimensional parameter space $u$ (for the emitted photon) – $a_{0}$ (external field intensity) – $p\cdot k/m^{2}$ (energy), where the constant crossed field approach is trustable. Since the differential cross section matches inevitably the Klein-Nishina cross section at certain (small) values of $u$, the total cross section contains also such a contribution. We argue that in discussing the Ritus-Narozhny conjecture one should resort to such a differential consideration. It could be that the tails of the Klein-Nishina behavior require adequate treatment in simulation codes, far beyond the improved constant cross field approximations.
For significant external field intensities, $a_{0}>1$, the Klein-Nishina effect, i.e. the suppression of (differential) cross section at high energies, becomes noticeably stronger.
Among the directions of further improvements are extensions accounting for polarization effects, e.g. along the path paved in Chen:2022dgo with various applications Seipt:2023bcw ; Gong:2022qis ; Song:2024oak ; Titov:2024aez . Much more compatibility (both computationally and w.r.t. parameter settings) is to be faced for nonlinear Compton scattering within a background field, which we envisage in subsequent work.
Appendix A A few Numbers
To justify the use of the considered fairly wide spread of energies and field strengths we quote a few numbers relevant to high-intensity laser physics and astrophysics.
$\lambda\ [\mathrm{nm}]$ $\omega\ [\mathrm{eV}]$ $I\ [\frac{\mathrm{W}}{\mathrm{cm}^{2}}]$ $a_{0}$101010 Using $a_{0}\approx 60\frac{\lambda}{\mu\mathrm{m}}\sqrt{I/(10^{22}\mathrm{W/cm}{}^{2})}$ with laser intensity $I$ and wave length $\lambda$. $\ E\ [\mathrm{GeV}]$ $\ pk/m^{2}$ $\ \ \chi$ ref.
LUXE 800 1.55 $2\times 10^{21}$ 21.5 17.5 0.2077 4.5 Abramowicz:2021zja
FACET-II
E320 800 1.55 $2\times 10^{20}$ 6.8 13 0.1543 1.05 Meuren:2020nbw
HED-HIBEF [^11] The primary aim is at vacuum birefringence investigations by two-laser interactions Ahmadiniaz:2024xob with (i) XFEL: $\omega=1\cdots 15$ keV, $a_{0}<\mathcal{O}(10^{-3})$ and (ii) optical laser: $\omega=1.55$ eV, $a_{0}<30$. These may optionally be used in collisions with (a) electron beam $E=100$ MeV (projection) or (b) rest gas electrons $E=\mathcal{O}(1~{}\mathrm{MeV})$ which are useful for calibration purposes in stand-alone configurations uncovering a wide parameter range, pointing either to the onset of the KN suppression regime, e.g. in (i.a), or to the suppression by the laser field in the Thomson regime, e.g. in (ii,a).
EuXFEL $0.1$ $13\times 10^{3}$ $10^{21}$ $0.002$ $0.1$ $9.5$ $0.02$ EuXFEL
$0.001$ $0.09$ $1.6\times 10^{-4}$
RELAX $795$ $1.56$ $8\times 10^{21}$ $42.7$ $0.1$ 0.001 $0.05$ RELAX
$0.001$ $1.1\times 10^{-5}$ $4.7\times 10^{-4}$
ELI-NP $814$ $1.52$ $10^{22}$ $49.8$ Gales:2018
CoReLS [^12] Current maximum for optical lasers. $800$ $1.55$ $1.1\times 10^{23}$ $159.2$ Yoon:2021ony
cosmic rays $10^{6}$ $10^{4}$ $7.7\times 10^{7}$ Boncioli:2023gbl
vela pulsar $10^{13}$ $2\times 10^{4}$ $1.5\times 10^{15}$ HESS:2023sxo
CMB $2.34\times 10^{-4}$ $10^{4}$ $0.0179$ ParticleDataGroup:2022pth
In general, magnetars produce magnetic fields131313 Critical field strengths: $B_{c}=m^{2}c^{3}/\hbar e\approx 4.4\times 10^{13}$ G and $E_{c}=m^{2}c^{3}/\hbar e\approx 1.323\times 10^{18}$ V/m. with $B\approx\mathcal{O}(10^{14}\cdots 10^{15}~{}\mathrm{G})$ Li:2024oda .
Appendix B Linear vs. circular polarization and CCF
[…]
lin. pol. circ. pol. CCF
$F_{1}=\mathcal{C}\omega~{}\times$ $\cos k\cdot x$ $\cos k\cdot x$ 1
$F_{2}=\mathcal{C}\omega~{}\times$ 00 $-\sin k\cdot x$ 0
$E_{x}=B_{y}=\mathcal{C}\omega~{}\times$ $\cos\Omega\tau$ $\cos\Omega\tau$ 1
$E_{y}=B_{x}=\mathcal{C}\omega~{}\times$ 00 $\sin\Omega\tau$ 0
$E=B=\mathcal{C}\omega~{}\times$ $\cos\Omega\tau$ 1 1
$T_{\mu\nu}=\mathcal{C}^{2}k_{\mu}k_{\nu}~{}\times$ $\cos^{2}k\cdot x$ 1 1
$a_{0}^{2}=\frac{e^{2}}{m^{2}}\mathcal{C}^{2}~{}\times$ $\frac{1}{2}$ 1 1
$b_{1}=\mathcal{C}~{}\times$ $1-\cos\Omega\tau$ $1-\cos\Omega\tau$ $\frac{1}{2}\Omega^{2}\tau^{2}$
$\approx\frac{1}{2}\Omega^{2}\tau^{2}(1+\frac{1}{12}\Omega^{2}\tau^{2})$ $\approx\frac{1}{2}\Omega^{2}\tau^{2}(1+\frac{1}{12}\Omega^{2}\tau^{2})$
$b_{2}=\mathcal{C}~{}\times$ 00 $-(\Omega\tau-\sin\Omega\tau)$ 0
$\approx-\frac{1}{6}\Omega^{3}\tau^{3}(1-\frac{1}{20}\Omega^{2}\tau^{2})$
$b_{3}=\mathcal{C}^{2}~{}\times$ $\frac{1}{4}(2\Omega\tau-\sin 2\Omega\tau)$ $2(\Omega\tau-\sin\Omega\tau)$ $\frac{1}{3}\Omega^{3}\tau^{3}$
$\approx\frac{1}{3}\Omega^{3}\tau^{3}(1-\frac{1}{5}\Omega^{2}\tau^{2})$ $\approx\frac{1}{3}\Omega^{3}\tau^{3}(1-\frac{1}{20}\Omega^{2}\tau^{2})$
where $\Omega\tau:=k\cdot x$ and $k=\omega n$. (We use coordinates with $n=(1,0,0,1)$, $\epsilon_{1}=(0,1,0,0)$, $\epsilon_{2}=(0,0,1,0)$, hence $E_{z}=B_{z}=0$.) The equality of $T_{\mu\nu}$, $E$ and $B$, and $a_{0}^{2}$ for circular polarization and CCF supports their close correspondence. The electron trajectories determine the classical emission. They are described by $x^{\mu}(\tau)=x^{\mu}(0)+p^{\mu}(0)\,\frac{\tau}{m}-\epsilon^{\mu\,i}b_{i}(\tau)\frac{e}{p\cdot k}+k^{\mu}b_{3}(\tau)\frac{e^{2}}{2(p\cdot k)^{2}}$ with coefficients $b_{1,2,3}$ listed above. In leading order in powers of $\Omega\tau$, all $b_{1,3}$ agree and $b_{2}^{\mathrm{circ.~{}pol.}}$ is nonzero but suppressed, $b_{2}^{\mathrm{circ.~{}pol.}}\approx-\frac{1}{3}b_{1}^{\mathrm{circ.~{}pol.}}\Omega\tau(1+\frac{1}{30}\Omega^{2}\tau^{2})$. The periodicity of the IPA trajectories (figure-8/ellipse in $x$-$z$ plane and a frame, where drift terms vanish, for lin./circ. polarizations) is at the origin of the IPA quasi-momenta, see Appendix in Seipt:2012tn for instance. Note the cancellation of $p\cdot k$ in the definition of $a_{0}^{2}$ and $a_{0}=\frac{e}{m}\mathcal{C}=\frac{m}{\omega}\frac{E}{E_{c}}$ due to $\mathcal{C}=E/\omega$ for both circular polarization and CCF. As emphasized above, $E$ and $\omega$ are primarily given by the considered plane wave and used afterward to define the related $a_{0}$ for CCF. Alternatively, one simply follows Ilderton:2018nws and defines local $\chi(\phi)$ and $a_{0}(\phi)$, thus considering CCF expressions as some approximation of IPA without recourse to real CCF.
Appendix C Master curves of harmonic structures
For $u\ll 1$ and $a_{0}\gg 1$, the leading order of $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{IPA}}}{du}$ depends only on $\mathfrak{u}:=u/u_{1}\approx ua_{0}^{2}m^{2}/2p\cdot k$ since $z_{n}\approx 2\sqrt{\mathfrak{u}(n-\mathfrak{u})}+\mathcal{O}(a_{0}^{-2})$. Harmonic structures occupy the region $\mathfrak{u}=\mathcal{O}(1)$. That is, $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{IPA}}}{du}$ serves as a unique master curve irrespective of the actual values of $u\ll 1$ or $p\cdot k/m^{2}$ or $a_{0}\ll 1$. For instance, keeping $p\cdot k/m^{2}$ and $a_{0}$ constant generates the master curve in Fig. 6-left, $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{IPA}}}{du}$ vs. $\mathfrak{u}=u/u_{1}$, as a concise representation of harmonic structures around $u/u_{1}\approx 1$, as exhibited in Fig. 2 as a function of $u$. Keeping $u$ and $a_{0}$ constant, the master curve in Fig. 6-right for the harmonic structures in Fig. 3 as a function of $p\cdot k/m^{2}$ is generated; here $\mathfrak{u}^{-1}=\frac{p\cdot k}{m^{2}}\frac{2}{ua_{0}^{2}}$ serves a bottom axis. The extension to a master curve as a function of $a_{0}$ scaled by $\sqrt{um^{2}/2p\cdot k}$ is obvious. The representation as a function of $a_{0}^{2}\frac{um^{2}}{p\cdot k}$ reproduces Fig. 6-left.
Figure 6: Master curves $u_{1}\frac{d\sigma^{\mathrm{IPA}}}{du}$ displaying harmonic structures on a severe background at $u\ll 1$ and $a_{0}\gg 1$ as a function of $u/u_{1}=\mathfrak{u}$ (left panel, complementing Fig. 2) and $u_{1}/u=\frac{p\cdot k}{m^{2}}\frac{2}{u(1+a_{0}^{2})}=\mathfrak{u}^{-1}$ (right panel, complementing Fig. 3). One can identify five harmonics $n=1,\cdots,5$, with the last one ($n=5$) being very shallow. Due to independent evaluations (the left panel is for running $u$ with $p\cdot k/m^{2}=10$ and $a_{0}=10^{2}$, while the right panel is for running $p\cdot k/m^{2}$ with $u=10^{-2}$ and $a_{0}=10^{2}$), the curves are not perfectly mirror-symmetric, as they should be in the asymptotic region. The l.h.s. triangle is for $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{KN}}}{du}|_{u\to 0}$, and the r.h.s. bullet is at $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u/u_{1}=10}$, both in the left panel only.
When $a_{0}<1$, the second and higher harmonics do not form local maxima; instead, the $u$ spectra become down stair-like. (Pulses with envelope function $g(\phi)=1/\cosh(\phi/10\pi)$ exhibit a similar – but further smoothed – spectrum, cf. Fig. 2 in Acosta:2021iyu .) Analogously, larger values of $u$, $u>1$, facilitate structureless distributions as a function of $a_{0}$, displaying the “rise and fall” behavior as in Fig. 3 in Acosta:2021iyu , where also pulses are considered for $a_{0}\leq 1$. Reference King:2020hsk emphasizes the appearance of an additional mid-IR peak, related to interference and $n=0$ contributions in pulses for $a_{0}=\mathcal{O}(1)$.
Appendix D $\omega^{\prime}$ and $\Theta^{\prime}$ differential distributions
After the concise consideration of the invariant differential cross section $d\sigma^{\mathrm{IPA}}/du$ and its dependencies, we turn to the differential cross section $d\sigma^{\mathrm{IPA}}/d\omega^{\prime}$, where $\omega^{\prime}=\nu^{\prime}m$ is the energy (scaled frequency) of the $out$ photon. One cannot directly employ $d\sigma^{\mathrm{IPA}}/du$, Eq. (9), to get $d\sigma^{\mathrm{IPA}}/d\omega^{\prime}$, instead one must use a differently weighted summation, in general,
$\frac{d\sigma^{\mathrm{IPA}}}{d\omega^{\prime}}=-\frac{2\pi\alpha^{2}}{p\cdot k}\sum_{n=1}^{N\to\infty}\Theta(u_{n}-u(n,\nu^{\prime}))\frac{1}{m\kappa_{n}}\mathcal{F}_{n}(\hat{z}_{n}),$ (22)
where $\hat{z}_{n}$ refers to the definition Eq. (11) with $u\to u(n,\nu^{\prime}):=\frac{n\nu-\nu^{\prime}}{\kappa_{n}-n\nu+\nu^{\prime}}$, see Kampfer:2020cbx ; Harvey:2009ry for details. The quantity $\kappa_{n}:=n\nu-\frac{1}{2}e^{\zeta}+\frac{1}{2}(1+a_{0}^{2})e^{-\zeta}$ with electron rapidity $\zeta=\mathrm{Arcosh}E_{e^{-}}/m$ can be positive or negative, meaning read or blue shift. Here, $\nu:=\omega/m$.
Figure 7: The quantity $u(n=1,\Theta^{\prime})$ as a function of the photon scattering angle $\Theta^{\prime}$ for various values of $a_{0}=1,\cdots,10^{4}$. The r.h.s. endpoints are at $u(n=1,\Theta^{\prime}=\pi)=u_{1}$ (cf. Eq. (13)). To accomplish larger values of $u$ at given $\Theta^{\prime}$, $n>1$ is required, pointing towards the CCF regime. In contrast, in the region where $u(n=1,\Theta^{\prime})\approx u^{\mathrm{KN}}(\Theta^{\prime})$, the differential cross section is saturated by the first harmonic (cf. Fig. 2) and $d\sigma^{\mathrm{IPA}}/du\approx d\sigma^{\mathrm{KN}}/du$ applies (cf. Eq. (18)). The chosen value $p\cdot k/m^{2}=0.1$ arises from $e^{\zeta}=5\times 10^{4}$ and $\nu=\omega/m=2\times 10^{-6}$. Note the scaling $u(n,\Theta^{\prime})=nu(1,\Theta^{\prime})$. The faint squares are for the KN kinematics $u^{\mathrm{KN}}(\Theta^{\prime}):=\frac{2\nu(1-\cos\Theta^{\prime})}{e^{\zeta}(1+\cos\Theta^{\prime})+e^{-\zeta}(1-\cos\Theta^{\prime})}$.
The differential cross section $d\sigma^{\mathrm{IPA}}/d\omega^{\prime}$ depends now on four parameters: $\omega^{\prime}$ (for the emitted photon in the lab.) and $a_{0}$ (external field intensity) and $\omega$ (frequency of the external field in lab.) and $E_{e^{-}}$ (electron initial energy in lab.), while the invariant differential cross section $d\sigma^{\mathrm{IPA}}/du$ had only three parameters, $u$ and $a_{0}$ and $p\cdot k/m^{2}$, with the latter one being a combination of $\omega$ and $E_{e^{-}}$ for head-on collisions. That is why the invariant differential cross section allows for a better survey of dependencies. For similarities and differences of both cross sections and peculiarities w.r.t. red shift vs. blue shifts, cf.  Harvey:2009ry . In an experiment, where both the frequency $\omega^{\prime}$ and angle $\theta^{\prime}$ of the $out$ photon are individually registered in each scattering event, the variable $u$ can be constructed by $u=\frac{e^{-\zeta}\nu^{\prime}(1-\cos\theta^{\prime})}{1-e^{-\zeta}\nu^{\prime}(1-\cos\theta^{\prime})}$. Without fixing the respective harmonic number $n$, the inversions $u\mapsto\omega^{\prime}$ or $u\mapsto\theta^{\prime}$ are not possible, in general. Considering $u(n,\Theta^{\prime})=\frac{2n\nu e^{\zeta}(1-\cos\Theta^{\prime})}{e^{2\zeta}(1+\cos\Theta^{\prime})+(1+a_{0}^{2})(1-\cos\Theta^{\prime})}$ Harvey:2009ry , instead of $u(n,\omega^{\prime})$, yields the pattern exhibited in Fig. 7. Of course, no dynamics is included here; it is simply kinematics aimed at exposing the statements in Ilderton:2018nws ; Blackburn:2021rqm : the small-$u$ region refers to large scattering angles too. For the considered kinematics, $u$ of the first harmonic, $n=1$, is independent of the intensity parameter $a_{0}$ as long as $a_{0}\leq 10^{4}$ and $\Theta^{\prime}<0.8\pi$ (see inclined arrow); for smaller $a_{0}$ this region extends towards the backscattering region. In the displayed example, $\mathfrak{s}r_{\perp}^{2}\approx 0.2$, where $r_{\perp}:=\frac{1}{\mathfrak{s}}\nu^{\prime}\sin\Theta^{\prime}$ is the scaled transverse momentum of the $out$ photon and, again, $\mathfrak{s}=\frac{u}{1+u}$. Such relations are at work w.r.t. the coincidence of IPA with KN in the small-$u$ region, see Fig. 2, also showing that restricting simulation codes to on-axis back scattering-only (vertical arrow) is not perfect, and beaming Blackburn:2019lgk must be included.
Acknowledgements.
The authors gratefully acknowledge the former collaboration with D. Seipt, T. Nousch, T. Heinzl, A. Otto and useful discussions with A. Ilderton, K. Krajewska, M. Marklund, C. Müller, S. Rykovanov, and G. Torgrimsson. A. Ringwald and B. King are thanked for explanations w.r.t. LUXE. The work is supported by R. Sauerbrey and T. E. Cowan w.r.t. the study of fundamental QED processes for HIBEF. The present documentation (being a survey from a particle physics perspective, thus our preference of cross sections) has been initiated by the HZDR HIBEF group. Useful discussions with R. Schützhold, C. Kohlfürst, and N. Ahmadiniaz are thanked for. The work of UHA was partly funded by the Center for Advanced Systems Understanding (CASUS) that is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament.

### v3

@@
 In doing so we resort to fairly known formalism as a vehicle to elucidate, numerically, links between established descriptions of linear Compton scattering, nonlinear Compton scattering, and photon emission in a constant crossed field.
+In this context, we not only review well-known results but also extend the comparative analysis of these findings, which are, in part, widely dispersed across highly specialized publications.
 This subject is also of interest for approximations implemented in simulation codes - both for astrophysics applications and laser physics in particular (cf.
@@
 It is known LLIV ; Ritus85 to emerge from nonlinear Compton scattering, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, as leading-order weak-field term, where $e_{A}^{-}$ is for the electron wave function in an external, classical background field $A$ which may refer to a monochromatic laser field in a plane-wave model as the so-called infinite plane-wave approximation (IPA).
-Having this relation in mind, one can define a suitably normalized probability (henceforth dubbed ”cross section”) which makes this relation explicit:444 See Bragin:2020akq for the definition of “local cross sections” for processes in an external field.
+Having this relation in mind, one can define a suitably normalized probability [^4] The relation of Ritus’ differential probability Ritus85 and the differential cross section is given by Ivanov:2004fi ; Acosta:2021iyu $\mathrm{d}W/\mathrm{d}u=a_{0}^{2}m^{2}(p\cdot k)/(e^{2}q^{0})\,\mathrm{d}\sigma/\mathrm{d}u$, where $q^{0}$ is the zeroth component of the effective momentum of the incoming electron.
+(henceforth dubbed ”cross section”) which makes this relation explicit:555 See Bragin:2020akq for the definition of “local cross sections” for processes in an external field.
 Instead of using wave packages, we employ a circularly polarized, monochromatic plane-wave field as a proxy of a mono-energetic photon beam under special circumstances.
@@
 (9), is often required.
-(Based on properties of the Bessel functions of large arguments and large indices, harmonics up $N=\mathcal{O}(a_{0}^{3})$ should be accounted for.) In accomplishing that goal, we employ the series expansion method referred to as Richardson extrapolation Sidi:2003bnm , which drastically improves the rate of conversions.
+(Based on properties of the Bessel functions of large arguments and large indices, harmonics up $N=\mathcal{O}(a_{0}^{3})$ should be accounted for.) In accomplishing that goal, we employ the series expansion method referred to as Richardson extrapolation Sidi:2003bnm , which drastically improves the rate of convergence.
 $\sigma^{\mathrm{IPA}}$ is compared with $\sigma^{\mathrm{KN}}$ in Fig.
@@
 As far as we know, this knowledge appears to be relatively uncommon.
-Sometimes, it is claimed in the literature that large values of $\chi$ would refer to high intensities, and the differential probability behind Eq.
+At large values of $a_{0}$, the differential probability behind Eq.
 (9) turns in the expression for one-photon emission in a constant crossed field (CCF).
@@
 Turning now to the energy dependence we exhibit in Fig.
-3 the differential IPA cross section for $u=0.1$ (left panel), $u=1$ (middle panel) and $u=10$ (right panel).
+3 the differential IPA cross section for $u=0.1$ (top panel), $u=1$ (left bottom panel) and $u=10$ (right bottom panel).
 For a given value of $a_{0}$, the IPA cross section matches the CCF expression at small energies, while the high-energy tails coincide with KN.
@@
 To identify further regions of dominating CCF behavior and discriminate against the KN regime, we display in Fig.
-4 the differential cross section at $u=0.1$ (right panel), $u=1$ (middle panel) and $u=10$ (right panel) as contour plots over the $a_{0}$-$p\cdot k/m^{2}$ plane.
+4 the differential cross section at $u=0.1$ (top panel), $u=1$ (left bottom panel) and $u=10$ (right bottom panel) as contour plots over the $a_{0}$-$p\cdot k/m^{2}$ plane.
 The differential cross section clearly shows that even for such large values of $\chi$ as $10^{4}$ one encounters the KN regime at large energy $p\cdot k/m^{2}>10^{2}$.
@@
 The $a_{0}$-$p\cdot k/m^{2}$ plane can be divided in three regions, YKNexp.supp.CCF at $u=const$, all captured by $d\sigma^{\mathrm{IPA}}/du$ at once.
-Such a survey might help discuss applicability domains of LCFA and LMA which however require the $u$ integration to arrive at total rates, as pointed out in footnote 6.
-Our consideration suggest that, on a differential level, LMA turns also into LCFA at high intensities.
-VI Many roads lead to large $\chi$
+Similarly, this was envisaged in Khokonov:2005 , but is here both quantified and expressed differentially in $u$.
+Such a survey might help discuss the applicability domains of LCFA and LMA which however require the $u$ integration to arrive at total rates, as pointed out in footnote 7.
+Our consideration suggests that, on a differential level, LMA turns also into LCFA at high intensities.
+VI Accuracy analysis for IPA vs CCF
+Figure 5: Example of absolute (left) and relative error (right) as a function of $u$ for $p\cdot k/m^{2}=0.1$ and $a_{0}=10$.
+Besides being lengthy examined in previous work DiPiazza:2018bfu ; DiPiazza:2017raw ; Ilderton:2018nws ; Heinzl:2020ynb , the validity of the replacement $\mathrm{IPA}\to\mathrm{CCF}$ for given triple $(a_{0},\chi,u)$ can be controlled for the differential cross-section by the condition
+$\frac{u}{N_{\mathrm{ctrl}}}\geq\frac{2\chi}{a_{0}(1+a_{0}^{2})}=\frac{p\cdot k}{m^{2}}\frac{2}{1+a_{0}^{2}},$ (22)
+where $N_{\mathrm{ctrl}}$ is the number of harmonics included in IPA (cf.
+Eq.
+(9)).
+Technically, the usage of CCF (instead of IPA), e.g.
+in a simulation code, for each triple $(a_{0},\chi,u)$ where (22) is fulfilled, introduces an approximation error controlled by $N_{\mathrm{ctrl}}$.
+In Fig.
+5-left, the absolute approximation error $\mathrm{abs.err.}:=\left|d\sigma^{\mathrm{IPA}}/du-\sigma^{\mathrm{CCF}}/du\right|$ is exemplarily exhibited.
+It decreases for $u\geq u_{\mathrm{th}}=\frac{p\cdot k}{m^{2}}\frac{2N_{\mathrm{ctrl}}}{1+a_{0}^{2}}$.
+This approves again the statement that CCF becomes increasingly better for larger $u$.
+Accordingly, in Fig.
+5-right, this is quantified using the relative approximation error $\mathrm{rel.err.}=\frac{\mathrm{abs.err.}}{d\sigma^{\mathrm{IPA}}/du}$.
+For instance, including only $10$ harmonics, i.e., $N=10$, in the IPA (see l.h.s.
+of the respective dashed curve) and using CCF on the r.h.s.
+of this dashed curve introduces a maximum error in the single-digit-percent range and improves [^10] The increasing relative error in Fig.
+5-right for $u>1$ has no physical meaning and results from the imprecise numerical cancelations.
+again for larger $u$.
+Consequently, for a given $N_{\mathrm{ctrl}}>5$, the maximum approximation error is given as the value at the respective lower boundary for $u$, i.e.
+the error value for $(a_{0},\chi,u)$ which solved the equality of (22).
+Fig.
+6-left shows, that, for $a_{0}>1$, the maximum relative approximation error decreases with increasing $N_{\mathrm{ctrl}}$.
+Furthermore, it is revealed that the error for a given $N_{\mathrm{ctrl}}>5$ does not change drastically, besides numerical cancelations, for a wide range of $a_{0}>1$, which shows the universality of the control parameter $N_{\mathrm{ctrl}}$.
+This can also be observed in the contour plots in Fig.
+4, where a straight line gives the lower boundary of the CCF area in the transition to the IPA.
+As already discussed in Ilderton:2018nws , in the realm of the local constant-field approximation, the transition from IPA and CCF can be derived analytically by expanding $\mathrm{d}\sigma^{\mathrm{IPA}}/\mathrm{d}u$ in the parameter $\lambda:=\left(\frac{8pk}{m^{2}a_{0}^{2}u}\right)^{\frac{1}{3}}$ and only keep terms linear in $\lambda$.
+Then, the replacement of IPA with CCF is valid [^11] Similarly, in DiPiazza:2018bfu ; DiPiazza:2017raw , the parameter $\eta_{\mathrm{LCFA}}:=\frac{\chi}{a_{0}^{3}u}$ is proposed to check the validity of LCFA, which is applicable if $\eta_{\mathrm{LCFA}}\ll 1$.
+In this regard, using condition (22) leads to $1\gg 1/N_{\mathrm{ctrl}}\geq 2\chi/(a_{0}(1+a_{0}^{2})u)\approx\chi/(a_{0}^{3}u)=\eta_{\mathrm{LCFA}}$ for $N_{\mathrm{ctrl}}>1$ and large $a_{0}\gg 1$.
+if $\lambda\ll 1$.
+Using the solution of the equality of the condition (22), the expansion parameter $\lambda$ can be numerically related to the maximum relative error.
+In Fig.
+6-right, we observe that even if $\lambda$ is relatively large, say $\sim 0.7$, the introduced maximum error by using CCF instead of IPA is still in the single-digit percent range.
+Therefore, by using the condition (22) and especially the control parameter $N_{\mathrm{ctrl}}$, one can quantify, what the $\ll$” in $\lambda\ll 1$ actually means.
+Figure 6: Example of the maximum relative error as a function of $N_{\mathrm{ctrl}}$ (left) and $\lambda$ (right) for various values of $a_{0}=1,10,100,1000$, and $p\cdot k/m^{2}=0.1$
+VII Many roads lead to large $\chi$
 The $a_{0}$-$p\cdot k/m^{2}$ plane can be mapped out by many other coordinates, e.g.
@@
 An example of orthogonal coordinates is the pair $\chi,\eta^{2}$ with $\eta^{2}(a_{0},p\cdot k/m^{2})=(p\cdot k/m^{2})^{2}-a_{0}^{2}$ for $\eta^{2}\in[-\infty,\infty]$ and our $\chi(a_{0},p\cdot k/m^{2})=a_{0}\,p\cdot k/m^{2}$ for $\chi\in[0,\infty]$.
-Trading CCF as proxy of $\frac{d\sigma^{IPA}}{du}$, the differential cross section can be fitted by $\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u=const}=\mathcal{A}(u)\,\chi^{-x}$ for $u\in[0.1,20]$ in the asymptotic region, i.e.
+Trading CCF as proxy of $\frac{d\sigma^{\mathrm{IPA}}}{du}$, the differential cross section can be fitted by $\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u=const}=\mathcal{A}(u)\,\chi^{-x}$ for $u\in[0.1,20]$ in the asymptotic region, i.e.
 at large $\chi$.
@@
 $p\cdot k/m^{2}$ plane are exhibited in Fig.
-5-left, for definiteness starting at $a_{0}=1$ and $p\cdot k/m^{2}=1$.
+7-left, for definiteness starting at $a_{0}=1$ and $p\cdot k/m^{2}=1$.
 We find for a path along $a_{0}=const$ (black line): $x=1/3$, along $a_{0}=(p\cdot k/m^{2})^{\beta}$: $x=2/3$ for $\beta=1/2$ (red line), $x=5/6$ for $\beta=1$ (blue line), $x=1$ for $\beta=2$ (green line), and along $p\cdot k/m^{2}=const$ (magenta line): $x=4/3$.
@@
 (See remark in figure caption on the relation of $\eta^{2}\lessgtr 0$ sectors.) As said above, many ways lead to large $\chi$ (and result in fairly different asymptotic regimes).
+[…]
 Figure 7: Left panel: The $a_{0}$-$p\cdot k/m^{2}$ plane is mapped out by curves $\chi=const$ (grey lines).
@@
 In that region, where CCF is a good approximation of IPA, this finding applies equally well to $d\sigma^{\mathrm{IPA}}/du|_{u=const}$.
-VII Discussion of dependencies and regimes
+VIII Discussion of dependencies and regimes
 The here employed formulas (1 - 16) of one-photon emission in some external univariate null-fields display often structures such as $1+u$, $1+a_{0}^{2}$, $1-u/u_{n}$, or $1+u_{\mathrm{KN}}$ etc.
@@
 In other words, the three-dimensional parameter space has a layer from $u=F^{-1}\,\frac{2}{u}\frac{1}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}$ to $u=F\,\frac{2}{u}\frac{1}{1+a_{0}^{2}}\frac{p\cdot k}{m^{2}}$; above (below) that layer, CCF (KN) approximates fairly well the invariant differential IPA cross section.
-It comes as some surprise that $d\sigma^{\mathrm{IPA}}_{\mathrm{circ.\,pol.}}\to d\sigma^{\mathrm{CCF}}$, since circular polarization (circ.
+Although, it can be analytically proven Ritus85 , it is somewhat surprising that $d\sigma^{\mathrm{IPA}}_{\mathrm{circ.\,pol.}}\to d\sigma^{\mathrm{CCF}}$, since circular polarization (circ.
 pol.) means a rotating electric field, i.e.
@@
 pol.) wave fields.
-In Appendix B, this issue is considered, as also the meaning of $a_{0}$ w.r.t.
+In Appendix B, this issue is examined, along with a discussion of the meaning of $a_{0}$ w.r.t.
 CCF.
@@
 In momentum balance equations, the electron IPA quasi-momenta are replaced by momenta, as occurring in CCF and KN, of course.
-VIII Summary
+IX Summary
 Inspired by the ongoing discussion of the Ritus-Narozhny conjecture concerning the validity of the Furry picture for processes at and in external fields, we analyze numerical probabilities of linear and nonlinear Compton scattering.
 To establish a link to the Klein-Nishina cross section for linear Compton scattering, we chose suitably normalized probabilities as “cross sections”, mainly based on the long-known relation of the weak-field limit of the nonlinear Compton scattering matching the Klein-Nishina expression.
-Choosing furthermore the monochromatic plane wave model as a proxy of a mono-energetic photon beam, we show with established expressions [^9] Since we envisage a clear cut picture with well-defined precise basics, we resort to the simplest (quantum) formulas describing exclusive one-photon emission.
+Choosing furthermore the monochromatic plane wave model as a proxy of a mono-energetic photon beam, we show with established expressions [^12] Since we envisage a clear-cut picture with well-defined precise basics, we resort to the simplest (quantum) formulas describing exclusive one-photon emission.
 Thus, such a compendium ignores much of the refinements elaborated during the last years, e.g.
 w.r.t.
-pulse shapes (implying bandwidth, ponderomotive and interference effects), laser focusing, and the many useful approximation schemes and simulation codes etc., including the efforts of benchmarks.
+pulse shapes (implying bandwidth, ponderomotive and interference effects), laser focusing, and the many useful approximation schemes and simulation codes, etc., including the efforts of benchmarks.
 By exhibiting a few relevant plots, we want to display systematically in an intriguing manner the interplay of Klein-Nishina (i.e.
@@
 To justify the use of the considered fairly wide spread of energies and field strengths we quote a few numbers relevant to high-intensity laser physics and astrophysics.
-$\lambda\ [\mathrm{nm}]$ $\omega\ [\mathrm{eV}]$ $I\ [\frac{\mathrm{W}}{\mathrm{cm}^{2}}]$ $a_{0}$101010 Using $a_{0}\approx 60\frac{\lambda}{\mu\mathrm{m}}\sqrt{I/(10^{22}\mathrm{W/cm}{}^{2})}$ with laser intensity $I$ and wave length $\lambda$.
+$\lambda\ [\mathrm{nm}]$ $\omega\ [\mathrm{eV}]$ $I\ [\frac{\mathrm{W}}{\mathrm{cm}^{2}}]$ $a_{0}$131313 Using $a_{0}\approx 60\frac{\lambda}{\mu\mathrm{m}}\sqrt{I/(10^{22}\mathrm{W/cm}{}^{2})}$ with laser intensity $I$ and wave length $\lambda$.
 $\ E\ [\mathrm{GeV}]$ $\ pk/m^{2}$ $\ \ \chi$ ref.
@@
 CMB $2.34\times 10^{-4}$ $10^{4}$ $0.0179$ ParticleDataGroup:2022pth
-In general, magnetars produce magnetic fields131313 Critical field strengths: $B_{c}=m^{2}c^{3}/\hbar e\approx 4.4\times 10^{13}$ G and $E_{c}=m^{2}c^{3}/\hbar e\approx 1.323\times 10^{18}$ V/m.
+In general, magnetars produce magnetic fields161616 Critical field strengths: $B_{c}=m^{2}c^{3}/\hbar e\approx 4.4\times 10^{13}$ G and $E_{c}=m^{2}c^{3}/\hbar e\approx 1.323\times 10^{18}$ V/m.
 with $B\approx\mathcal{O}(10^{14}\cdots 10^{15}~{}\mathrm{G})$ Li:2024oda .
+In addition to the listed laser facilities, we mention NSF ZEUS ZEUS , NSF OPAL OPAL , and SEL SELabc .
+For further details, the interested reader is referred to the cited references.
 Appendix B Linear vs. circular polarization and CCF
@@
 For instance, keeping $p\cdot k/m^{2}$ and $a_{0}$ constant generates the master curve in Fig.
-6-left, $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{IPA}}}{du}$ vs.
+8-left, $\frac{p\cdot k}{m^{2}}\frac{d\sigma^{\mathrm{IPA}}}{du}$ vs.
 $\mathfrak{u}=u/u_{1}$, as a concise representation of harmonic structures around $u/u_{1}\approx 1$, as exhibited in Fig.
@@
 Keeping $u$ and $a_{0}$ constant, the master curve in Fig.
-6-right for the harmonic structures in Fig.
+8-right for the harmonic structures in Fig.
 3 as a function of $p\cdot k/m^{2}$ is generated; here $\mathfrak{u}^{-1}=\frac{p\cdot k}{m^{2}}\frac{2}{ua_{0}^{2}}$ serves a bottom axis.
@@
 The representation as a function of $a_{0}^{2}\frac{um^{2}}{p\cdot k}$ reproduces Fig.
-6-left.
+8-left.
 Figure 8: Master curves $u_{1}\frac{d\sigma^{\mathrm{IPA}}}{du}$ displaying harmonic structures on a severe background at $u\ll 1$ and $a_{0}\gg 1$ as a function of $u/u_{1}=\mathfrak{u}$ (left panel, complementing Fig.
@@
 Considering $u(n,\Theta^{\prime})=\frac{2n\nu e^{\zeta}(1-\cos\Theta^{\prime})}{e^{2\zeta}(1+\cos\Theta^{\prime})+(1+a_{0}^{2})(1-\cos\Theta^{\prime})}$ Harvey:2009ry , instead of $u(n,\omega^{\prime})$, yields the pattern exhibited in Fig.
-7.
+9.
 Of course, no dynamics is included here; it is simply kinematics aimed at exposing the statements in Ilderton:2018nws ; Blackburn:2021rqm : the small-$u$ region refers to large scattering angles too.
@@
 The work of UHA was partly funded by the Center for Advanced Systems Understanding (CASUS) that is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament.
+Data Availability Statement
+This manuscript has no associated data. The code used to produce the plots in this manuscript are available upon request from the corresponding author.
+Author Contribution Statement
+All authors contributed equally to the preparation of this manuscript.